# Meta-generalized-gradient exchange-correlation functionals for spin-orbit coupled electrons

### v1

@@ Abstract
The prominence of density functional theory (DFT) in the field of electronic structure computation stems from its ability to usefully balance accuracy and computational effort. At the base of this ability is a functional of the electron density: the exchange-correlation (xc) energy. This functional satisfies known exact conditions that guide the derivation of approximations. The strongly-constrained-appropriately-normed (SCAN) approximation stands out as a successful, modern, example. In the presence of spin-orbit coupling, however, the xc energy must depend not only on the electron density but also on spin currents – a fact which is not captured by SCAN or by any mainstream functional approximation. In this work, we demonstrate how the SU(2) gauge-invariance of the exact functional allows us to add the necessary dependence on spin currents in the SCAN functional (here called JSCAN) — and similar meta-generalized-gradient functional approximations — in a non-empirical manner. The prescription further ensures that the conjugate xc-fields of the spin currents have vanishing local torques as in the exact case. In passing, a spin-current dependent generalization of the electron localization function (here called JELF) is also derived. The extended forms are implemented in a developer’s version of the Crystal23 program. Applications on molecules and materials confirm the practical relevance of the extensions.
@@
71.15.Mb, 71.15Rf, 31.15.E-Introduction. Relativistic effects play an important role in our understanding of the electronic structure of matter. Greiner (2012) Spin-orbit coupling (SOC), in particular, emerges naturally in the two-component Pauli equation (a simplified low-energy version of the Dirac equation) and plays a crucial role in our understanding of the band structure of materials such as Si, Ge and GaAs, which powered the semiconductor revolution. Yu and Cardona (2010) More recently, the interest in spin-orbit interactions in condensed matter has grown even further following the discovery of a host of materials which exhibit nontrivial topological properties Bernevig and Hughes (2013); Vanderbilt (2018) and the explosion of interest in new techniques to accurately control the spin of the electrons in spintronic and quantum computing devices.utić et al. (2004)
Just as an electric potential couples to the particle density and a magnetic field couples to the spin and to the particle (i.e. ordinary) current, SOC can be described as a tensor field which couples to the spin current (SC), i.e., the tensor product of spin and ordinary current. Unlike the ordinary current, the spin current is invariant under time reversal (being the product of two time-reversal odd quantities) and therefore can freely flow in the time-reversal invariant ground state of a non magnetic system.Tokatly (2008); Droghetti et al. (2022)
@@
However, there is no reason why a density functional approximation that was developed for systems lacking SOC would still yield accurate results when strong SOC are present. To accomplish this more ambitious task, we need an exchange-correlation energy functional that depends not only on the density but also — and crucially — on the spin current density, and does so in a universal (i.e., system-independent) manner, as mandated by the original spin-current density functional theory (SCDFT)Vignale and Rasolt (1988); Bencheikh (2003) and its more recent developments. Rohra and Görling (2006); Heaton-Burgess et al. (2007); Trushin and Görling (2018); Desmarais et al. (2023a) In practice, this means including in the effective potential, not only the “bare” SOC, but also its “dressing” due to many-body effects, in analogy to the dressing of the bare electron-nuclear potential by Hartree and exchange-correlation (xc) effects.
@@
But DFT (and thus SCAN) cannot deal with non-perturbative SOC! This brings forth the necessity to switch to SCDFT both in principleVignale and Rasolt (1988); Bencheikh (2003) and in practice.Rohra and Görling (2006); Heaton-Burgess et al. (2007); Trushin and Görling (2018); Desmarais et al. (2023a, 2020a); Bodo et al. (2022); Comaskey et al. (2022) The dependence of the xc energy functional on spin currents in SCDFT is strongly constrained by the requirement that its form remains invariant under local SU(2) gauge transformations (defined below). Taking as a starting point the SCAN and the ELF forms, in this work we show how to make them functionals of the spin currents while preserving the fundamental SU(2) gauge invariance of the theory. The resulting generalizations are here called JELF and JSCAN. A set of computed results for molecules and materials demonstrates the practical usefulness of our non-empirical extension of the SCAN functional. An implementation is presented in a developer’s version of the Crystal23 program.Erba et al. (2023)
SCDFT in a nutshell. We shall consider ground states at geometries that do not magnetize nor host particle currents spontaneously. Furthermore, magnetic couplings (Zeeman field, Abelian vector potential) that break time-reversal symmetry are not included. Accounting for SOC, however, we must consider together with the particle density, , the spin currents, , and corresponding external couplings: a scalar, , and a non-Abelian vector potential, . The former accounts for the electron-nuclear potential, while the latter provides a useful formal representation of SOC.Fröhlich and Studer (1993); Bencheikh (2003); Abedinpour et al. (2010); Zhang and Vignale (2016); Droghetti et al. (2022) The energy density functional of interest, thus, reads as follows:
@@
Invoking a MGGA, it is convenient to consider the so-called generalized-KS approach to SCDFT.Desmarais et al. (2023a) This means that for generating the single-particle equations, we differentiate the energy functional expression explicitly w.r.t. the orbitals rather than w.r.t. to the densities not (a) — nowadays, this has become a standard procedure in DFT.Neumann et al. (1996)
@@
It is important to notice that the spin current itself is not a gauge invariant quantity since it can appear as a consequence of a generic SU(2) transformation acting on a state in which . The invariance of stated in Eq. (3), implies that only certain gauge-invariant combinations of with other KS quantities can legitimately appear in . We now tackle the task of generalizing the SCAN functional (and the closely related ELF) to include a dependence on spin currents that complies with Eq. (3).
Extension of SCAN and ELF to spin-current carrying states. The SCAN functional setsSun et al. (2015) where interpolates between the semi-local energy densities for single-orbital densities () and for slowly-varying densities (), which only depend on (a gauge-invariant quantity). The interpolation function is controlled by the variable . The latter quantity entails three kinetic energy densities: the positive-definite conventional defined in terms of the occupied (one-component) KS orbitals {}, the von Weizsäcker — that is the bosonic expression of — where , and the kinetic energy of the noninteracting Fermi gas at uniform density : .
@@
Note, the above substitution should not be confused with the analogous well-known substitution that involves the (paramagnetic) particle current:Tao and Perdew (2005) The latter derives from the consideration of local U(1) transformations. The importance of the dependence of density functionals on the particle current has been largely demonstrated. Dobson (1993); Becke (1996b, 2002); Maximoff et al. (2004); Burnus et al. (2005); Tao and Perdew (2005); Pittalis et al. (2007, 2009); Räsänen et al. (2009); Oliveira et al. (2010); Zhu et al. (2016); James W. Furness and Teale (2016); Holzer et al. (2022)
Next, from Eq. (5), taking a functional derivative with respect to , we obtain
@@
For the non-magnetic ground states considered in this work, the JSCAN satisfies the non-Abelian torque conditions exactly. not (c)
@@
Applications. We have implemented a self-consistent treatment of the JSCAN functional, including effective non-Abelian vector potentials, as well as the JELF in a developer’s version of the Crystal23 code.Erba et al. (2023) Computational and implementation details are provided in the supporting informationESI (see also Refs. Desmarais et al., 2021a, 2020b, 2019a, 2019b; Metz et al., 2000; Peterson et al., 2003, 2007; Doll et al., 2001; Doll, 2001; Doll et al., 2006; Civalleri et al., 2001; Laun et al., 2018; Heyd et al., 2005; Lebedev, 1976, 1977; Towler et al., 1996; Lehtola et al., 2020 therein). In the calculations, we employ the restored-regularized r-SCAN, rather than the original SCAN functional, to improve numerical stability and performance.Furness et al. (2022)
Exx-Only | Exx & Corr | |||
J-rSCAN | HF | J-rSCAN | RPAx | |
FI | 0.248 (0.225) | 0.322 | 0.232 (0.222) | 0.360 |
HI | 0.337 (0.276) | 0.343 | 0.324 (0.281) | 0.343 |
I | 0.333 (0.276) | 0.364 | 0.317 (0.278) | 0.396 |
@@
We apply our SU(2) gauge-invariant J-rSCAN approach to molecules on reproducing the SOC-induced modification of the valence electronic structure in halogen diatoms and hydrides, which have served as model systems in previous studies.Desmarais et al. (2021b, c, 2023b) Values of the SOC-induced modification to the HOMO-LUMO gap are reported in Table 1, first at the exchange-only level of theory, and then including effects of exchange and correlation. Values are presented with exchange-only (gauge-invariant) J-rSCAN functional, and numbers in parentheses represent corresponding results with the standard (gauge-dependent) rSCAN approach.
@@
Next, we discuss electron localization through the JELF [see Eq. (6)]. Let us begin by recalling that large values of , corresponding to a (J)ELF close to one, indicate that electrons are highly localized. Here we look at the effect of SOC on the (J)ELF. For the I molecule, after a self-consistent solution with J-rSCAN, including both exchange and correlation effects, in Fig. 1(a) we plot the ELF and JELF indicators as differences w.r.t. the calculation without SOC. Yellow contours indicate that SOC localizes the electron. Blue contours indicate that the electron is delocalized by SOC. The leftmost panel, containing the JELF shows that electrons are localized along the bond axis by SOC. On the other hand, we observe two blue “onion rings” around the atomic centers. Thus, SOC localizes electrons along the bond and delocalizes them in the orthogonal directions. Although this analysis is consistent with previous reports,Desmarais et al. (2023c); Gulans and Draxl (2022) the degree of delocalization is largely overestimated by the gauge-dependent ELF, compared to the SU(2) form-invariant JELF, see their difference in the rightmost panel of Fig. 1(a). The JELF displays more localization on the bonding compared to the (gauge-dependent) ELF.
@@
We proceed to apply our approach to the description of the electronic band structure of infinite periodic systems. In Table 2, we provide values of the splitting of the valence band at the high-symmetry point in molybdenum dichalcogenide Rashba systems: the inversion asymmetric hexagonal single layer MoSe and the inversion-symmetric hexagonal bulk -MoTe. The valence band structure is provided in panel (b) of Figure 1.
@@
For both systems, about one third of the total effect of SOC on the splitting is accounted for through many-body effects by inclusion of spin currents in the functional. Low-temperature experimental values on the transport gap are available from Refs. Choi et al., 2017; Island et al., 2016; Zelewski and Kudrawiec, 2017; Conan et al., 1979, while band splittings are provided in Refs. Oliva et al., 2020; Zhang et al., 2014; Shim et al., 2014; Ross et al., 2013; Beal et al., 1972; Ruppert et al., 2014. Including spin currents in the xc form has a very small effect on the band gaps (see Table 2). On the other hand, the values of valence band splittings are considerably enhanced (from 0.14 eV to 0.20 eV in MoSe, and from 0.28 eV to 0.32 eV in MoTe), leading to an improved agreement with experiment (relative error decreased from 29% to 10% in MoSe and leading to an exact coincidence with the mean experimental value in MoTe). Overall, the accuracy of the J-rSCAN approach is comparable to more involved hybrid-GGA SCDFT treatments (which provide splittings of 0.17 and 0.32 eV, respectively for MoSe and MoTe using optimal fractions of Fock exchange) — without requiring the determination of an optimal fraction of Fock exchange:Desmarais et al. (2023a) a step that involves either empiricism or extra/external ab initio calculations.Skone et al. (2014)
2D MoSe | 3D -MoTe | |||
split | gap | split | gap | |
rSCAN | 0.14 | 1.57 | 0.28 | 0.82 |
J-rSCAN | 0.20 | 1.53 | 0.32 | 0.81 |
@@
Finally, we report on the application of our SU(2) form-invariant approach to topological materials. We consider the case of Weyl semi-metals. We report on the splitting of Weyl nodes in the orthorhombic TaAs phase. We study the W1 node pair, which lies along the mirror plane. The pair of nodes is located at (in units of ), in excellent agreement with the experimental value of .Lv et al. (2015); Huang et al. (2015) Spin-current dependent terms in the xc functional account for of the total splitting along . See Fig. S1 of the supplementary material.
Conclusions. Focusing on the prominent example of the SCAN functional, we have shown how to include a dependence on spin currents in a meta-GGA xc functional form for evaluating SOC effects in such a way as to fulfill the prerequisite of SU(2) gauge invariance of the xc energy. For the states considered in this work, the current-dependent form of the approximation is furthermore strongly suggested by previous studies on the structure of the exact exchange hole (see Ref. Pittalis et al., 2017), which incidentally, also motivate the inclusion of the kinetic energy density as a variable on which the functional must depend. A large body of works have highlighted the usefulness of the kinetic energy in combination with other quantities as in the ELF (and, thus, as in the iso-orbital indicator of SCAN) as a practical means for capturing relevant local features of the many-electron state in correlated regimes beyond exchange-only effects. Of course, the deeper question of including the “right physics” beyond JSCAN remains open. Aside from exact constraints provided by the SU(2) gauge principle, the success of the benchmark calculations reported in our paper leads us to believe that we are proceeding on the right path.
Looking ahead, the approach proposed in this work will help disentangle the role of the SU(2) invariance in density functional approximations from questions more directly related to the physics of the electron-electron interaction. It is particularly appealing to explore the consequence of a full (compound) U(1)SU(2) gauge-invariance on general states that may also break time-reversal symmetry. This may help solve standing difficulties of present DFT methodologies for magnetized systems.Isaacs and Wolverton (2018); Jana et al. (2018); Ekholm et al. (2018); Fu and Singh (2018, 2019); Mejía-Rodríguez and Trickey (2019); Romero and Verstraete (2018); Buchelnikov et al. (2019); Tran et al. (2020)

### v2

# Spin-currents via the gauge-principle for meta-generalized-gradient exchange-correlation functionals

@@ Abstract
The prominence of density functional theory (DFT) in the field of electronic structure computation stems from its ability to usefully balance accuracy and computational effort. At the base of this ability is a functional of the electron density: the exchange-correlation energy. This functional satisfies known exact conditions that guide the derivation of approximations. The strongly-constrained-appropriately-normed (SCAN) approximation stands out as a successful, modern, example. In this work, we demonstrate how the SU(2) gauge-invariance of the exchange-correlation functional in spin current density functional theory allows us to add an explicit dependence on spin currents in the SCAN functional (here called JSCAN) — and similar meta-generalized-gradient functional approximations — solely invoking first principles. In passing, a spin-current dependent generalization of the electron localization function (here called JELF) is also derived. The extended forms are implemented in a developer’s version of the Crystal23 program. Applications on molecules and materials confirm the practical relevance of the extensions.
@@
71.15.Mb, 71.15Rf, 31.15.E-Introduction. Spin current carrying states have acquired a major role in condensed matter physics and electronic device engineering. Greiner (2012) Recently, the interest in spin currents has grown further following the discovery of a host of materials which exhibit nontrivial topological properties Bernevig and Hughes (2013); Vanderbilt (2018) and the development of new techniques to accurately control the spin of the electrons in spintronic and quantum computing devices.utić et al. (2004)
Commonly, spin currents originate from spin-orbit coupling (SOC). In many situations, SOC can be accounted for via the two-component Pauli equation, which offers formal and computational simplification as compared to the full Dirac equation. Just as an electric potential couples to the particle density and a magnetic field couples to the spin and to the particle (i.e. ordinary) current, SOC can be described as a tensor field which couples to the spin current (SC), i.e., the tensor product of spin and ordinary current. Unlike the ordinary current, the spin current is invariant under time reversal (being the product of two time-reversal odd quantities) and therefore can freely flow in the time-reversal invariant ground state of a non magnetic system.Tokatly (2008); Droghetti et al. (2022)
@@
However, there is no reason why a pure density functional approximation that was developed for systems lacking spin currents would still yield accurate results when SOC is present. To accomplish this more ambitious task — yet within the two-component Pauli equation — we need an exchange-correlation energy functional that depends not only on the density but also — and crucially — on the spin current density, and does so in a universal (i.e., system-independent) manner, as mandated by the original spin-current density functional theory (SCDFT)Vignale and Rasolt (1988); Bencheikh (2003) and its more recent developments. Rohra and Görling (2006); Heaton-Burgess et al. (2007); Trushin and Görling (2018); Desmarais et al. (2024) In practice, this means including in the effective potential, not only the “bare” SOC, but also its “dressing” due to many-body effects, in analogy to the dressing of the bare electron-nuclear potential by Hartree and exchange-correlation (xc) effects.
@@
But regular DFT cannot deal with spin-currents induced by extra interactions. This brings forth the necessity to switch to SCDFT both in principleVignale and Rasolt (1988); Bencheikh (2003) and in practice.Rohra and Görling (2006); Heaton-Burgess et al. (2007); Trushin and Görling (2018); Desmarais et al. (2024); Desmarais et al. (2020a); Bodo et al. (2022); Comaskey et al. (2022) The dependence of the xc energy functional on spin currents in SCDFT is strongly constrained by the requirement that its form remains invariant under local SU(2) gauge transformations (defined below). Taking as a starting point the SCAN and the ELF forms, in this work we show how to make them functionals of the spin currents while preserving the fundamental SU(2) gauge invariance of the theory. The resulting generalizations are here called JELF and JSCAN. A set of computed results for molecules and materials demonstrates the practical usefulness of our non-empirical extension of the SCAN functional. An implementation is presented in a developer’s version of the Crystal23 program.Erba et al. (2023)
SCDFT in a nutshell. We shall consider ground states at geometries that do not magnetize nor host particle currents spontaneously. Furthermore, magnetic couplings (Zeeman field, Abelian vector potential) that break time-reversal symmetry are not included. Accounting for non-Abelian vector potentials, however, we must consider together with the particle density, , the spin currents, .Fröhlich and Studer (1993); Bencheikh (2003); Abedinpour et al. (2010); Zhang and Vignale (2016); Droghetti et al. (2022) The energy density functional of interest, thus, reads as follows:
@@
Invoking a MGGA, it is convenient to switch to the so-called generalized-KS (GKS) approach to SCDFT.Desmarais et al. (2024) This means that for generating the single-particle equations, we differentiate the energy functional expression explicitly w.r.t. the orbitals rather than w.r.t. to the densities not (a) — nowadays, this has become a standard procedure in DFT.Neumann et al. (1996)
@@
It is important to notice that is not a gauge invariant quantity since it can appear as a consequence of a generic SU(2) transformation acting on a state in which . The invariance of stated in Eq. (3), implies that only certain gauge-invariant combinations of with other KS quantities can legitimately appear in . We now tackle the task of generalizing the SCAN functional (and the closely related ELF) to include a dependence on spin currents that complies with Eq. (3).
Extension of SCAN and ELF to spin-current carrying states. Recent works have moved valuable steps in developing extended DFT approaches to non-collinear magnetism and SOC.Pittalis et al. (2017); Holzer et al. (2022); Tancogne-Dejean et al. (2023); Desmarais et al. (2024) But Ref. Pittalis et al., 2017 does not carry out applications, Ref. Holzer et al., 2022 does not work with full SU(2)-invariant xc-functionals, Ref. Tancogne-Dejean et al., 2023 does not include an xc-vector potential in the solution, and Ref. Desmarais et al., 2024 does not include spin-current explicitly beyond exchange.
Here, we focus on MGGAs. The SCAN functional, for example, setsSun et al. (2015) where interpolates between the semi-local energy densities for single-orbital densities () and for slowly-varying densities (), which only depend on (a gauge-invariant quantity). The interpolation function is controlled by the variable . The latter quantity entails three kinetic energy densities: the positive-definite conventional defined in terms of the occupied (one-component) KS orbitals {}, the von Weizsäcker — that is the bosonic expression of — where , and the kinetic energy of the noninteracting Fermi gas at uniform density : .
@@
Note that, under the restriction to time-reversal symmetric states, the rest of the SCAN only depends on the particle density (a fully invariant quantity) and, thus, shall not be modified here. Also note that , as for the original quantity. Importantly, the above substitution should not be confused with the analogous well-known substitution that involves the (paramagnetic) particle current:Tao and Perdew (2005) The latter derives from the consideration of local U(1) transformations. The importance of the dependence of xc functionals on the particle current has been already largely demonstrated. Dobson (1993); Becke (1996b, 2002); Maximoff et al. (2004); Burnus et al. (2005); Tao and Perdew (2005); Pittalis et al. (2007, 2009); Räsänen et al. (2009); Oliveira et al. (2010); Zhu et al. (2016); James W. Furness and Teale (2016)
For the states considered in this work as for any gauge-invariant MGGA within GKS, we stress that does not transform covariantly and does not exert a torque on the spin current. In fact, from Eq. (5), taking a functional derivative with respect to , we obtain
@@
Yet, the role of is far from negligible (more below). The GKS equations for the JSCAN are reported in the supporting informationESI in a form which is handy for numerical implementations.
@@
Applications. We have implemented a self-consistent treatment of the JSCAN functional, including effective non-Abelian vector potentials, as well as the JELF in a developer’s version of the Crystal23 code.Erba et al. (2023) Computational and implementation details are provided in the supporting informationESI (see also Refs. Desmarais et al., 2021a, 2020b, 2019a, 2019b; Metz et al., 2000; Peterson et al., 2003, 2007; Doll et al., 2001; Doll, 2001; Doll et al., 2006; Civalleri et al., 2001; Laun et al., 2018; Heyd et al., 2005; Lebedev, 1976, 1977; Towler et al., 1996; Lehtola et al., 2020 therein). In the calculations, we employ the restored-regularized r2-SCAN, rather than the original SCAN functional, to improve numerical stability and performance.Furness et al. (2022)
Exx-Only | Exx & Corr | ||||
J-r2SCAN | HF | J-r2SCAN | RPAx | ||
FI | 0.248 (0.225) | 0.322 | 0.232 (0.222) | 0.360 | |
HI | 0.337 (0.276) | 0.343 | 0.324 (0.281) | 0.343 | |
I2 | 0.333 (0.276) | 0.364 | 0.317 (0.278) | 0.396 | |
@@
We apply our SU(2) gauge-invariant J-r2SCAN approach to molecules on reproducing the SOC-induced modification of the valence electronic structure in halogen diatoms and hydrides, which have served as model systems in previous studies.Desmarais et al. (2021b, c, 2023a) Values of the SOC-induced modification to the HOMO-LUMO gap are reported in Table 1, first at the exchange-only level of theory, and then including effects of exchange and correlation. Values are presented with exchange-only (gauge-invariant) J-r2SCAN functional, and numbers in parentheses represent corresponding results with the standard (gauge-dependent) r2SCAN approach.
@@
Next, we discuss electron localization through the JELF [see Eq. (6)]. We recall that large values of , corresponding to a (J)ELF close to one, indicate that electrons are highly localized. Here we look at the effect of SOC on the (J)ELF. For the I2 molecule, after a self-consistent solution with J-r2SCAN, including both exchange and correlation effects, in Fig. 1(a) we plot the ELF and JELF indicators as differences w.r.t. the calculation without SOC.
@@
Yellow contours indicate that SOC localizes the electron. Blue contours indicate that the electron is delocalized by SOC. The leftmost panel, containing the JELF shows that electrons are localized along the bond axis by SOC. On the other hand, we observe two blue “onion rings” around the atomic centers. Thus, SOC localizes electrons along the bond and delocalizes them in the orthogonal directions. Although this analysis is consistent with previous reports,Desmarais et al. (2023b); Gulans and Draxl (2022) the degree of delocalization is largely overestimated by the gauge-dependent ELF, compared to the SU(2) form-invariant JELF, see their difference in the rightmost panel of Fig. 1(a). The JELF displays more localization on the bonding compared to the (gauge-dependent) ELF.
@@
We proceed to apply our approach to the description of the electronic band structure of infinite periodic systems. In Table 2, we provide values of the splitting of the valence band at the high-symmetry point in molybdenum dichalcogenide Rashba systems: the inversion asymmetric hexagonal single layer MoSe2 and the inversion-symmetric hexagonal bulk -MoTe2. The valence band structure is provided in panel (b) of Figure 1.
@@
For both systems, about one third of the total effect of SOC on the splitting is accounted for through many-body effects by inclusion of spin currents in the functional. Low-temperature experimental values on the transport gap are available from Refs. Choi et al., 2017; Island et al., 2016; Zelewski and Kudrawiec, 2017; Conan et al., 1979, while band splittings are provided in Refs. Oliva et al., 2020; Zhang et al., 2014; Shim et al., 2014; Ross et al., 2013; Beal et al., 1972; Ruppert et al., 2014. The values of valence band splittings are considerably enhanced by including spin-current in the xc form (from 0.14 eV to 0.20 eV in MoSe2, and from 0.28 eV to 0.32 eV in MoTe2), leading to an improved agreement with experiment (relative error decreased from 29% to 10% in MoSe2 and leading to an exact coincidence with the mean experimental value in MoTe2). Overall, the accuracy of the J-r2SCAN approach is comparable to more involved hybrid-GGA SCDFT treatments (which provide splittings of 0.17 and 0.32 eV, respectively for MoSe2 and MoTe2 using optimal fractions of Fock exchange) — without requiring the determination of an optimal fraction of Fock exchange:Desmarais et al. (2024) a step that involves either empiricism or extra/external ab initio calculations.Skone et al. (2014)
2D MoSe2 | 3D -MoTe2 | ||||
split | gap | split | gap | ||
r2SCAN | 0.14 | 1.57 | 0.28 | 0.82 | |
J-r2SCAN | 0.20 | 1.53 | 0.32 | 0.81 | |
@@
Finally, we report on the application of our SU(2) form-invariant approach to topological materials. We consider the case of Weyl semi-metals. We report on the splitting of Weyl nodes in the orthorhombic TaAs phase (W1 node pair, along the mirror plane). The pair of nodes is located at (in units of ), in excellent agreement with the experimental value of .Lv et al. (2015); Huang et al. (2015) Spin-current dependent terms in the xc functional account for of the total splitting along . See Fig. S1 of the supplementary material.
Conclusions. Focusing on the prominent example of the SCAN functional and the ELF, we have shown how to include a dependence on spin currents in meta-GGA functional forms while fulfilling SU(2) gauge invariance for time-reversal symmetric systems. For the states considered in this work, the current-dependent form of the approximation is furthermore strongly suggested by previous studies on the structure of the exact exchange hole (see Ref. Pittalis et al., 2017), which incidentally, also motivate the inclusion of the kinetic energy density as a variable on which the functional must depend. A large body of works have highlighted the usefulness of the kinetic energy in combination with other quantities as in the ELF (and, thus, as in the iso-orbital indicator of SCAN) as a practical means for capturing relevant local features of the many-electron state in correlated regimes beyond exchange-only effects.
Of course, the question of including the “right physics” beyond the minimalistic yet necessary approach presented here remains open. Looking ahead, it will help disentangle the role of the SU(2) invariance in density functional approximations from questions more directly related to the physics of the electron-electron interaction. It is particularly appealing to explore the consequence of a full (compound) U(1)SU(2) gauge-invariance on general states that may also break time-reversal symmetry. This may help solve standing difficulties of present DFT methodologies for magnetized systems.Isaacs and Wolverton (2018); Jana et al. (2018); Ekholm et al. (2018); Fu and Singh (2018, 2019); Mejía-Rodríguez and Trickey (2019); Romero and Verstraete (2018); Buchelnikov et al. (2019); Tran et al. (2020)